a particle moves along a straight line such that at time t its displacement from a fixed point o on the line is 3t^2 -2 the velocity of the particle when t=2 is

Grade:11

1 Answers

Abhishek Sharma
10 years ago

ans - 12m/s

just differtiate 3t^2 -2 and put t=2

on differentiation we get - 6t

and velocity = 6*2 = 12 m/ s

Please approve